FMARS Mission Support Summary - Thursday, July 12th
This has been the second day of rotation 2 under Dr Zubrin`s command. Before
leaving, Dr Lee's team completed 2 simulated EVA's to Hayne's Ridge and to
Haughton Crater. No signs of hydrotermal activity at the first site but lithic
microorganisms at the second. One EVA lasted 4.5 hours. IVA officers
documented the whole procedure during both EVAs which will be a standard
procedure from now on as it provides very useful insights on the operation and
support of this activity. Before leaving the Hab, the first crew reported that
their house had become comfortable and safe needing some more work to become
fully operational. Crew 2 already performed a 2 hour EVA (including a 1 hr
preparation time and 5 min. pre-breath) the night they arrived and found
stomatolites and macroscopic fossiles.The plan for today was an extensive EVA
at Haynes's ridge and a voice report was forwarded to Mission Support at 6 pm
with details to be further expanded in the science reports. The crew reported
to feeling very well. Food supply will arrive on Saturday with the help of
Aziz at Camp's Inn.
Mission support welcomed Katya Muscat as a new Capcom today. Keith Savage
worked on a providing the team with a daily satellite image and Brian Encke was
involved with the KSC communication protocol. At 5:00pm MDT Lorraine Bell took
over as Mission Director so that Tony Muscatello could take a much needed break.
There was a communication problem confirmed between crew 1 and Mission Support
on July 11th as messages have been forwarded to the FMARS address instead of
commander's personal email. Three messages were not received by Pascal Lee but
forwarded to him today : one about the Hull experiment, one informing that
Mission Support was going offline and about a request for a tentative schedule
to the next day's plan. The Hull experiment will be pursued with crewmembers'
The Capcom task list is being improved daily which will result in clearer
communication between ground and Field.